OpenStudy (wampominater):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (wampominater):

so far, i have followed the formula \[\cos(\theta) = \frac{ u*v }{\left| v \right| \left| u \right| }\] and i have gotten 1, but that isnt an answer choice...

OpenStudy (wampominater):

and then when I take the arc cos of 1, I get 0, which is not a choice

OpenStudy (anonymous):

\( \huge |u| = \sqrt{(-5)^2+(-4)^2} = \sqrt{41} \) \( \huge |v| = \sqrt{(-4)^2+(-3)^2} = 5 \) That what you have for |v| and |u| ?

OpenStudy (wampominater):

yeah, i believe i know what i did wrong now though

OpenStudy (anonymous):

you sure? I can work it out if you want

OpenStudy (wampominater):

well, i think that its because i rounded the denom to 32

OpenStudy (wampominater):

but sure, if you want to help me do it just to make sure, thatd be great! :)

OpenStudy (wampominater):

so i got approx 1.8 when i didnt round the denominator

OpenStudy (anonymous):

We should have \( \huge \cos(\theta) = \frac{ u*v }{\left| v \right| \left| u \right| }\) \( \huge \cos(\theta) = \frac{ 32}{\left| 5 \right| \left| \sqrt{41} \right| } = .99951 = 1.78 degrees\) That is what I got.

OpenStudy (wampominater):

awesome, that is what i got too, thank you for the help!

OpenStudy (anonymous):

YW!!!